# Edge-Only Universal Adversarial Attacks in Distributed Learning

### v1

@@ Abstract
Distributed learning frameworks, which partition neural network models across multiple computing nodes, enhance efficiency in collaborative edge-cloud systems but may also introduce new vulnerabilities. In this work, we explore the feasibility of generating universal adversarial attacks when an attacker has access to the edge part of the model only, which consists in the first network layers. Unlike traditional universal adversarial perturbations (UAPs) that require full model knowledge, our approach shows that adversaries can induce effective mispredictions in the unknown cloud part by leveraging key features on the edge side. Specifically, we train lightweight classifiers from intermediate features available at the edge, i.e., before the split point, and use them in a novel targeted optimization to craft effective UAPs. Our results on ImageNet demonstrate strong attack transferability to the unknown cloud part. Additionally, we analyze the capability of an attacker to achieve targeted adversarial effect with edge-only knowledge, revealing intriguing behaviors. By introducing the first adversarial attacks with edge-only knowledge in split inference, this work underscores the importance of addressing partial model access in adversarial robustness, encouraging further research in this area.
1 Introduction
In the last decade, adversarial attacks have posed significant challenges in deploying secure and robust deep learning models [7, 6, 34]. These attacks involve carefully crafted perturbations to input data that can effectively mislead model predictions, originating serious concerns about the safety and security of machine learning algorithms in various application domains, such as autonomous driving [30, 43] and healthcare [13].
@@ 1 Introduction
Figure 1: Scheme of the proposed edge-only universal attack. The attacker has full control only over the edge part, from which they can exploit the output of its layers to train binary classifiers $g^{l}$ that learn feature separations characterizing reference target samples. This is then exploited to run an adversarial optimization that crafts a universal attack transferable to the cloud part of the model.
@@
Contribution. To address these challenges, we propose a novel approach for generating universal adversarial attacks with knowledge limited to the initial layers of the model, which are commonly deployed at the edge side. The main idea, illustrated in Figure 1, relies on exploiting a small set of new samples to learn a binary separation of class-specific features from the available edge layers between a reference target class and all other classes. Even without access to the rest of the model, this separation can guide a targeted adversarial optimization to craft universal perturbations, steering attacked inputs towards stimulating certain features in the edge layers that lead to misclassification when propagated to the rest of the model.
We conducted several experiments on ImageNet [14], varying the number of layers deployed on the edge side and testing different models. Experimental results demonstrate that the proposed attack paradigm induces a high misclassification rate in the model output, even when only a small number of layers is accessible at the edge, achieving performance often comparable to white-box UAP attacks [40, 20], depending on the selected target class. Additionally, we provide an in-depth analysis of how targeted optimization propagates through the unknown cloud portion of the model, revealing that even when the attack is applied at the initial layers only, it shifts features towards specific unintended classes on the cloud side, effectively mimicking a targeted attack but towards an alternate class. In summary, this work makes the following contributions:
  • It introduces the first edge-only attack in distributed inference, where the cloud portion of the model is entirely unknown;
  • It presents multiple experiments comparing the effectiveness of edge-only UAPs with classic full-model knowledge attacks, across different target classes;
  • It reveals intriguing behaviors in the model outcomes with edge-only attacks, understanding the impact of targeting specific classes within the initial layers.
The remainder of the paper is organized as follows: Section 2 discusses the related work and the impact on the literature; Section 3 covers background terminology and formalizes UAPs; Section 4 details the methodology for crafting edge-only universal adversarial attacks; Section 5 presents experimental results on various split points of multiple DNN models; and finally, Section 6 provides conclusions, limitations, and directions for future work.
2 Related Work
@@
Universal Adversarial Attacks.
In recent years, adversarial perturbations have gained significant attention as a way to reveal critical weaknesses in the robustness of DNNs [34, 6, 17]. Among these, universal adversarial attacks are particularly important from a security standpoint, as they maintain their effectiveness across different inputs [32, 19, 20, 42]. Moosavi-Dezfooli et al.[20] were the first to demonstrate the existence of UAPs by introducing an adversarial optimization method that extends the DeepFool attack[21] to extract features that generalize across multiple images. Since then, various strategies have been developed to enhance UAPs’ effectiveness [24, 22]. In these studies, also targeted formulations of UAPs were explored, which are practically interesting as they avoid the need for labeled samples during optimization (i.e., the perturbation is optimized to push any input towards a specific class) [2, 20]. In addition, interesting studies have been conducted to understand the nature of UAPs, suggesting that they often relate to task-specific features rather than the training data itself [40, 41, 23, 25], and, when applied, these features become predominant across layers[40]. This property makes input-agnostic features, defined by the model’s weights, sufficient for crafting effective UAPs.
@@
Attacks and Threat Models in Distributed Learning.
The development of distributed and collaborative learning paradigms has enabled models to be partitioned across multiple nodes, enhancing computational efficiency in edge computing environments and IoT applications [18, 12]. However, this distributed nature also introduces new vulnerabilities and potential threat models [1]. Numerous studies have addressed issues related to privacy and authenticity in distributed systems, as in federated learning [16], backdoor attacks [35], and privacy attacks in split learning through manipulation of the cloud component [28, 9].
In contrast to previous studies, we explore the case of an attacker that could gain control over the edge part only at a post-training stage (distributed inference). This is a practical threat, as edge devices, often constrained by limited computational resources, may be more vulnerable to attacks, potentially allowing an attacker to cloning them or leveraging the gradients from the split point [8, 1]. In this context, no prior work has thoroughly examined adversarial optimization strategies that exploit knowledge of the edge component alone. Lacking direct access to the decision boundary of the model output, or the ability to estimate it through black-box techniques, we considered UAPs as useful tool for exploiting feature distributions only in preliminary layers.
3 Background and Modeling
In this section, we first recap key terminologies related to split inference in distributed learning, and the threat model we consider. We then formalize universal perturbations, with a specific focus on the targeted formulation that drives the approach proposed in Section 4.
Split Inference.
@@ 3 Background and Modeling
Threat Model.
We study a secure setting in which the model output is not returned to the edge part [28]. We assume that: (i) the attacker can collect a small set of samples that differ from the training set; (ii) the attacker has white-box access to the edge portion of the model only, i.e., $f_{\text{edge}}$, allowing to study its activations and perform gradient backpropagation; (iii) the complete set of output classes $\{1,\dots,N_{c}\}$ is unknown to the attackers, while they can focus on a specific class (target class) to analyze the feature space at the edge.
The practical relevance of these assumptions is motivated as follows: (i)-(ii) the attacker can gain access to the edge node, even by means of physical access, hence being able to leak both the edge portion of the model and input samples; (iii) the attacker can classify a small set of samples independently of the model $f$ under attack.
Universal Adversarial Perturbations
UAPs are particularly interesting from a security perspective due to their generalizability, as they maintain effectiveness across different samples without needing to be re-computed for each input. Formally, the objective is to find a perturbation $\delta$ that induces mispredictions on any input $x$ of a distribution $X$. This problem can be formulated as follows:
@@
where $y$ is the ground-truth output related to $x$, $\mathcal{L}$ is a loss function (e.g., cross-entropy loss), and $\epsilon$ is the maximum magnitude of $\delta$ under a $p$-norm constraint. In this work, we focus on the $l_{\infty}$ norm, commonly adopted in the UAP-related literature [41].
Particularly relevant to this work is the targeted variant of UAPs, where the goal is to induce predictions towards a specific target class $y_{t}$ selected by the attacker. Optimizing targeted attacks involves minimizing a loss function with respect to $y_{t}$:
@@
Let $D_{\text{opt}}$ be a dataset collected by the attacker that does not include inputs of the target class $y_{t}$ only, and without any particular restriction on its distribution. In practice, the perturbation $\delta$ can be iteratively computed by means of gradient descent:
@@
where $B$ represents a batch of samples from $D_{\text{opt}}$, $\alpha$ is the step size, and sign the $l_{\infty}$ projection. After each update a clip in $[-\epsilon,+\epsilon]$ is applied, to ensure the $\epsilon$ constraint.
4 Edge-Only Universal Attack
@@
In this section, we detail the edge-only universal attack following the pipeline shown in Figure 1. The attack consists of three main stages: (1) learning feature separations between target and non-target samples in the edge layers; (2) crafting the perturbation by exploiting these learned separations; and (3) applying an adversarial perturbation to influence the final model prediction.
4.1 Learning Class-Feature Separations
@@
where $\mathcal{L}_{b}$ denotes the binary cross-entropy loss, $\theta_{g^{l}}$ are the weights of $g^{l}$, and $\bar{y}_{t}$ is an auxiliary label set to 1 for target samples and to 0 for non-target samples. By this binary classifier the attacker learns which features most strongly represent the target class and use them to guide the optimization of targeted universal perturbations.
It is also important to note that the depth of the edge part of the model, defined as the number of layers $e$ deployed at the edge, plays a significant role in the generalization capability beyond the samples in $\mathcal{D}$. In particular, while knowledge of deeper layers improves the transferability of attacks to the cloud part, it also carries a higher risk of overfitting when considering few samples, reducing the ability to generalize. To address this problem, we found that learning feature separation from multiple edge layers and combining them in an optimization process is beneficial, as also shown in the ablation studies reported in Section 5.
4.2 Adversarial Optimization
In the adversarial optimization stage, we aim to craft a universal perturbation that pushes an arbitrary sample $x$ to be classified in the target class $y_{t}$ by the binary classifiers $\{g^{l_{1}},\dots,g^{l_{e}}\}$ introduced above, so that $g^{l}(f^{l}(x+\delta))=1$ for $l\in L_{\textit{edge}}$. This problem can be framed as a targeted universal optimization, where the objective is to craft malicious input patterns so that the perturbed features match those of the target class, thereby influencing the processing performed by the cloud part of the model. Extending the optimization problem of Eq. (2), the simplest approach to address this problem consists in adopting gradient ascent to each classifier $g^{l}$ as follows:
$\delta_{t+1}=\delta_{t}+\alpha\cdot\text{sign}\left(\sum_{l\in L_{\text{edge}}}\sum_{x\in B}\nabla_{\delta}g^{l}(f^{l}(x+\delta_{t}))\right)$ (5)
@@ 4.2 Adversarial Optimization
where each gradient $\nabla_{\delta}g^{l}(\cdot)$ of classifier $g^{l}$ is normalized by its $\ell_{2}$-norm to ensure balanced contributions across layers. Both the usage of multiple gradients and the normalization allow improving the transferability of the attack to the cloud part, as shown in studies reported in Section 5.4.
@@
$\delta_{t+1}=\delta_{t}+\alpha\cdot\text{sign}\left(\sum_{x\in B}\Phi(x,\delta_{t})\right)$ (7)
where $\|\delta\|_{\infty}\leq\epsilon$. The resulting perturbed samples can eventually be used to induce mispredictions, keeping their representation in $[0,1]$, i.e., $\tilde{x}=\textit{clip}(x+\delta,0,1)$.

### v2

@@ Abstract
Distributed learning frameworks, which partition neural network models across multiple computing nodes, enhance efficiency in collaborative edge-cloud systems, but may also introduce new vulnerabilities to evasion attacks, often in the form of adversarial perturbations. In this work, we present a new threat model that explores the feasibility of generating universal adversarial perturbations (UAPs) when the attacker has access only to the edge portion of the model, consisting of its initial network layers. Unlike traditional attacks that require full model knowledge, our approach shows that adversaries can induce effective mispredictions in the unknown cloud component by manipulating key feature representations at the edge. Following the proposed threat model, we introduce both edge-only untargeted and targeted formulations of UAPs designed to control intermediate features before the split point. Our results on ImageNet demonstrate strong attack transferability to the unknown cloud part, and we compare the proposed method with classical white-box and black-box techniques, highlighting its effectiveness. Additionally, we analyze the capability of an attacker to achieve targeted adversarial effects with edge-only knowledge, revealing intriguing behaviors across multiple networks. By introducing the first adversarial attacks with edge-only knowledge in split inference, this work underscores the importance of addressing partial model access in adversarial robustness, encouraging further research in this area.
I Introduction
In the last decade, adversarial attacks have posed significant challenges in deploying secure and robust deep learning models [1, 2, 3]. These attacks involve carefully crafted perturbations to input data that can effectively mislead model predictions, originating serious concerns about the safety and security of machine learning in various application domains, such as autonomous driving [4, 5] and healthcare [6].
@@ I Introduction
To address these challenges, this work formalizes a novel edge-only adversarial threat model and proposes methods for generating UAPs in both targeted and untargeted settings. For the targeted formulation, the main idea is to exploit a small auxiliary attack set to learn class-specific feature patterns from the accessible edge layers, by contrasting a reference target class against all other classes. For the untargeted variant, the objective is instead to induce deviations from clean feature representations. Despite the lack of access to the remaining model components, we show that edge features are sufficient to guide the optimization of universal perturbations, steering attacked inputs to cause misclassification when propagated through the unknown cloud-side model.
We conducted extensive experiments on ImageNet [26], varying both the number of layers deployed on the edge and the model architectures. The results show that the proposed attack paradigm induces high misclassification rates, even when only a few layers are accessible at the edge, achieving performance often comparable to white-box UAP methods [27, 9] in the untargeted setting. Comparable effectiveness is also observed for the targeted formulation, depending on the chosen target class and the depth of the accessible edge portion. Additionally, we provide an in-depth analysis of how targeted optimization propagates through the unknown cloud portion of the model, revealing that even when the attack is applied at the initial layers only, it shifts features towards specific unintended classes on the cloud side, effectively mimicking a targeted attack but towards a different class. In summary, this work makes the following contributions:
  • It introduces an edge-only threat model for distributed inference, where the cloud part of the model is unknown;
  • It presents formulations of both targeted and untargeted edge-only UAPs under the proposed threat model;
  • It reports extensive experimental results comparing the effectiveness of edge-only UAPs with traditional full-knowledge and black-box attacks, highlighting the practical relevance of addressing the proposed threat model.
The remainder of the paper is organized as follows: Sec. II discusses related work; Sec. III introduces background concepts; Sec. IV presents the threat model; Sec. V details the proposed methodologies for crafting edge-only universal adversarial attacks, first in the targeted and then in the untargeted setting; Sec. VI presents the experimental results; and finally, Sec. VII presents the conclusions and directions for future work [^1] Code available at https://github.com/retis-ai/EdgeUAP.
II Related Work
@@
II-A Universal Adversarial Attacks
In recent years, adversarial perturbations have gained significant attention as a way to reveal critical weaknesses in the robustness of DNNs [3, 2, 7]. Among these, universal adversarial attacks are particularly important from a security standpoint, as they maintain their effectiveness across different inputs [28, 29, 9, 8]. This enables attackers to generate the perturbation offline and apply it at runtime, making such attacks significantly more practical in real-world scenarios. Moosavi-Dezfooli et al.[9] were the first to demonstrate the existence of UAPs by introducing an adversarial optimization method that extends the DeepFool attack[30] to extract features that generalize across multiple images. Since then, various strategies have been developed to enhance UAPs’ effectiveness [31, 32]. In these studies, also targeted formulations of UAPs were explored, which are practically interesting as they avoid the need for labeled samples during optimization (i.e., the perturbation is optimized to push any input towards a specific class) [33, 9]. In addition, interesting studies have been conducted to understand the nature of UAPs, suggesting that they often relate to task-specific features rather than the training data itself [27, 34, 35, 36], and, when applied, these features become predominant across layers[27].
@@ II-A Universal Adversarial Attacks
II-B Threat Models in Distributed Learning
The development of distributed and collaborative learning paradigms has enabled the training and inference of AI models across multiple nodes, improving computational efficiency in edge computing environments and IoT applications [15, 16]. However, these distributed paradigms also introduces novel security vulnerabilities and broadens the threat landscape [17]. Prior research has explored various threats in distributed settings, including backdoor attacks in federated learning [37, 38] and privacy issues in split learning through manipulation of cloud-side components [22, 23].
From the perspective of evasion attacks, however, previous work has not yet explored scenarios beyond conventional adversarial settings, where attackers rely on access to model outputs, either through full model access in white-box attacks or in black-box attacks [39, 40, 41, 42]. To fill this gap, we investigate a previously underexplored but realistic threat scenario in distributed model partitioning: an attacker gaining control over the edge component during inference. As discussed in more detail in Sec.III, this threat model is particularly relevant due to the practical limitations of edge devices, which often operate with constrained computational resources and weak security measures. These characteristics make them vulnerable to model cloning, parameter extraction, and gradient analysis at the partitioning point [43, 17, 44].
To address this gap, we study how an attacker can exploit shallow representations using UAPs, leveraging only the feature distribution from the edge component without access to the full model or decision boundary. This approach enables new insights into the vulnerability of distributed inference under realistic deployment constraints.
III Background
This section recaps key terminologies related to split inference in distributed learning and universal perturbations.
III-A Split Inference.
@@ III Background
III-B Universal Adversarial Perturbations.
Formally, the objective is to find a perturbation $\delta$ that induces mispredictions on any input $x$ of a distribution $X$. This problem can be formulated as follows:
@@
where $y$ is the ground-truth output related to $x$, $\mathcal{L}$ is a loss function (e.g., cross-entropy loss), and $\epsilon$ is the magnitude of $\delta$ under a $p$-norm constraint. In this work, we focus on the $l_{\infty}$ norm, commonly adopted in the related literature [34].
Particularly relevant to this work is the targeted variant of UAPs, where the goal is to induce predictions towards a specific target class $y_{t}$ selected by the attackers, i.e., minimizing a loss function with respect to $y_{t}$:
@@
Let $D_{\text{opt}}$ be a dataset collected by the attacker that does not include inputs of the target class $y_{t}$ only, and without any particular restriction on its distribution. In practice, the perturbation $\delta$ can be iteratively computed by means of gradient:
@@
where $B$ represents a batch of samples from $D_{\text{opt}}$, $\alpha$ is the step size, and sign the $l_{\infty}$ projection. After each update a clip in $\epsilon,+\epsilon]$ is applied, to ensure the $\epsilon$ constraint.
IV Edge-Only Attack Threat Model
We study a secure setting in which the model output is not returned to the edge part [22]. It assumes that: (i) the attacker can collect a small set of samples that differ from the training set; (ii) the attacker has white-box access to the edge portion of the model only, i.e., $f_{\text{edge}}$, allowing to study its activations and perform gradient backpropagation; (iii) the complete set of output classes $\{1,\dots,N_{c}\}$ is unknown to the attackers, while they can focus on a specific class (the target class) that is assumed to be part of this set (but not strictly necessary). In this setting, the attacker’s goal is to craft an adversarial perturbation that causes the model (on the cloud side) to produce an incorrect prediction.
The practical relevance of the above assumptions is motivated as follows: (i)-(ii) the attacker can gain access to the edge node, even by means of physical access, hence being able to leak both the edge portion of the model and input samples; (iii) the attacker can classify a small set of samples independently of the model $f$ under attack.
IV-A Relevance of the Threat Model
The proposed threat model captures a setting in which the attacker has no access to model outputs and no knowledge of the full label space, but can exploit internal representations from an initial portion of the network (the edge part). This scenario is increasingly relevant due to the growing deployment of split and collaborative neural networks in sensor nodes and IoT systems [45, 16] (e.g., wearable health-monitoring devices, smart retail sensors), where resource-constrained edge devices perform early encoding for feature extraction, compression, and partial inference, while the remaining computation is offloaded to remote servers. For instance, representative examples are smart video surveillance and traffic monitoring, where edges perform on-device feature extraction and transmit intermediate embeddings to a cloud service for final recognition or analytics. In such systems, the edge device is physically exposed and can be accessed, probed, or tampered with by an adversary, while the cloud model and final predictions remain protected behind secure infrastructure. In these practical deployments, compromising the edge-side model or its activations is significantly more feasible than breaching the cloud backend [44, 24, 25], reflecting a realistic and underexplored attack surface that emerges in modern distributed inference pipelines.
Figure 1: Comparison between threat models. Left: the proposed edge-based attack, where the adversary has access to the edge portion of the split model. Right: a standard black-box setting, where the adversary relies on a surrogate model $s$ to craft transferable attacks.
IV-B Clarifications with respect to black-box attacks.
The presented threat model does not fall under the standard weak black-box setting. As illustrated in Fig.1, classical black-box attacks typically assume that the adversary has access to the model’s outputs (e.g., in the form of predicted labels or full confidence scores), while the internal architecture and parameters remain hidden. In such settings, it is common to assume that the attacker can train or acquire a surrogate model ($s$ in Fig.1), which is then used to craft adversarial examples that are expected to transfer to the target [32, 46, 47].
These assumptions, although widely adopted to evaluate transferability across architectures, are often optimistic in practice: in private and proprietary deployments, the attacker may lack access to sufficiently representative data or architectural priors to effectively train highly transferable surrogates.
In contrast, our work focuses on a fundamentally different scenario in which the input is processed locally at the edge and no prediction feedback is returned to the attacker. This setup cannot be reduced to a standard black-box model, as the attacker has access to internal parameters (up to the split point) but no access to output logits or labels.
V Edge-Only Universal Attack
Figure 2: Scheme of the proposed targeted edge-only universal attack (Secs. V-A and V-B). The attacker has full control only over the edge part, from which can exploit the output of its layers to train binary classifiers $g^{l}$ that learn feature separations characterizing reference target samples. This is then exploited to run an adversarial optimization that crafts a universal attack transferable to the cloud part of the model.
@@ V Edge-Only Universal Attack
In this section, we detail the proposed edge-only universal attacks. We first formulates a targeted attack, following the pipeline illustrated in Fig. 2. The method consists of two main stages: (i) learning feature-level separations between target and non-target samples in the edge layers (Sec. V-A); and (ii) crafting a universal perturbation by exploiting these learned separations (Sec. V-B). We then introduce, in Sec. V-C, an untargeted attack that causes misclassification on the cloud side without controlling the predicted class.
V-A Learning Class-Feature Separations for Targeted Attacks
@@
where $\mathcal{L}_{b}$ denotes the binary cross-entropy loss, $\theta_{g^{l}}$ are the weights of $g^{l}$, and $\bar{y}_{t}$ is an auxiliary label set to 1 for target samples and to 0 for non-target samples. By this binary classifier the attacker learns which features most strongly represent the target class and use them to guide the optimization.
It is also important to note that the depth of the edge part of the model, defined as the number of layers $e$ deployed at the edge, plays a significant role in the generalization capability beyond the samples in $\mathcal{D}$. In particular, while knowledge of deeper layers improves the transferability of attacks to the cloud part, it also carries a higher risk of overfitting when considering few samples, reducing the ability to generalize. To address this problem, we found that learning feature separation from multiple edge layers and combining them in an optimization process is beneficial, as also shown in the ablation studies reported in Sec. VI.
V-B Targeted Adversarial Attack
In the adversarial optimization stage, we aim to craft a universal perturbation that pushes an arbitrary sample $x$ to be classified in the target class $y_{t}$ by the binary classifiers $\{g^{l_{1}},\dots,g^{l_{e}}\}$ introduced above, so that $g^{l}(f^{l}(x+\delta))=1$ for $l\in L_{\textit{edge}}$. This problem can be framed as a targeted universal optimization, where the objective is to craft malicious input patterns so that the perturbed features match those of the target class, thereby influencing the processing performed by the cloud part of the model. Extending the optimization problem of Eq. (2), a direct approach to address this problem consists of adopting gradient ascent to each classifier $g^{l}$ as follows:
$\delta_{t+1}=\delta_{t}+\alpha\cdot\text{sign}\left(\sum_{l\in L_{\text{edge}}}\sum_{x\in B}\nabla_{\delta}g^{l}(f^{l}(x+\delta_{t}))\right),$ (5)
@@ V-B Targeted Adversarial Attack
where each gradient $\nabla_{\delta}g^{l}(\cdot)$ of classifier $g^{l}$ is normalized by its $\ell_{2}$-norm to ensure balanced contributions across layers. Both the usage of multiple gradients and the normalization allow improving the transferability of the attack to the cloud part, as shown in studies reported in Sec. VI-E.
@@
$\delta_{t+1}=\delta_{t}+\alpha\cdot\text{sign}\left(\sum_{x\in B}\Phi(x,\delta_{t})\right),$ (7)
where $\|\delta\|_{\infty}\leq\epsilon$ keeps the perturbation under the $\epsilon$-magnitude.
The resulting perturbed samples can eventually be used to induce mispredictions, keeping their representation in $[0,1]$, i.e., $\tilde{x}=\textit{clip}(x+\delta,0,1)$, assuming images are normalized between 0 and 1 for convenience.
As observed in the experimental part (Sec.VI), the feasibility of achieving high targeted misclassification in edge-only settings depends on the degree of separation of class feature representations available at the edge depth. This characteristic depends on the model architecture, the selected class, and the depth of the accessible edge layers, which may exhibit limited class separability. Nevertheless, we found that such perturbations can effectively influence the final predictions, leading to misclassification with respect to the correct label, i.e., $f(x+\delta)\neq y$, even when the perturbation does not consistently drive the prediction toward a specific target class. This observation motivates the definition of the next attack variant, which is specifically designed for untargeted objectives.
V-C Untargeted Adversarial Attack
We also introduce an untargeted adversarial formulation as a direct extension of the CosUAP method by Zhang et al. [27]. While CosUAP optimizes a universal perturbation by minimizing, in expectation, the cosine similarity between the output logits of clean and perturbed samples, we adapt this principle to the intermediate features extracted by the edge model and backpropagate gradients only through the white-box accessible layers. Formally, this can be expressed as:
$\min_{\|\delta\|_{p}\leq\epsilon,}\;\mathbb{E}_{x\sim X}\left[\mathcal{L}_{\text{cos}}(f_{\text{edge}}(x+\delta),f_{\text{edge}}(x))\right],$ (8)
where $\mathcal{L}_{\text{cos}}$ denotes the cosine similarity. In practice, $\delta$ is updated iteratively, as in the previous targeted case (Eq. 7), but without the need to train the auxiliary classifier $g$ here, since no class-specific feature learning is required. Following the approach in Eq.6, the update direction is defined here as:
$\Phi_{\text{unt}}(x,\delta)=-\sum_{l\in L_{\text{edge}}}\frac{\nabla_{\delta}\,\mathcal{L}_{\text{cos}}\!\left(f^{l}(x+\delta_{t}),f^{l}(x)\right)}{\left\|\nabla_{\delta}\,\mathcal{L}_{\text{cos}}\!\left(f^{l}(x+\delta_{t}),f^{l}(x)\right)\right\|_{2}}.$ (9)